### v1

@@ Abstract
Existing test-time adaptation (TTA) approaches often adapt models with the unlabeled testing data stream. A recent attempt relaxed the assumption by introducing limited human annotation, referred to as Human-In-the-Loop Test-Time Adaptation (HILTTA) in this study. The focus of existing HILTTA lies on selecting the most informative samples to label, a.k.a. active learning. In this work, we are motivated by a pitfall of TTA, i.e. sensitive to hyper-parameters, and propose to approach HILTTA by synergizing active learning and model selection. Specifically, we first select samples for human annotation (active learning) and then use the labeled data to select optimal hyper-parameters (model selection). A sample selection strategy is tailored for choosing samples by considering the balance between active learning and model selection purposes. We demonstrate on 4 TTA datasets that the proposed HILTTA approach is compatible with off-the-shelf TTA methods which outperform the state-of-the-art HILTTA methods and stream-based active learning methods. Importantly, our proposed method can always prevent choosing the worst hyper-parameters on all off-the-shelf TTA methods. The source code will be released upon publication.
@@ 1 Introduction
The disparity in distribution between training and testing data leads to poor generalization of deep neural networks. The data distribution of testing data is often unknown until the inference stage [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] and the distribution may experience constant shift [Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Yuan et al.(2023)Yuan, Xie, and Li, Press et al.(2023)Press, Schneider, Kümmerer, and Bethge, Lee and Chang(2024), Su et al.(2024a)Su, Xu, and Jia]. These practical challenges lead to the research into test-time adaptation (TTA) [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Su et al.(2022)Su, Xu, and Jia, Su et al.(2024a)Su, Xu, and Jia]. Despite achieving impressive results on a wide range of tasks, e.g. classification [Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt], image segmentation [Wang et al.(2023)Wang, Zhong, Wang, Chen, Ling, Wang, and Sebe] and object detection [Chen et al.(2023)Chen, Xu, Su, and Jia], the fully unsupervised TTA paradigm is still overwhelmed by remaining challenges including hyper-parameter tuning [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], continual distribution shift [Wang et al.(2022)Wang, Fink, Van Gool, and Dai], non-i.i.d. [Gong et al.(2022)Gong, Jeong, Kim, Kim, Shin, and Lee, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan, Yuan et al.(2023)Yuan, Xie, and Li], class imbalance [Su et al.(2024a)Su, Xu, and Jia], etc. Unfortunately, there are still no principled solutions to address these challenges without relaxing the assumptions.
Recently, efforts have been made to the exploration of introducing a limited annotation budget during test-time adaptation (SimATTA) [Gui et al.(2024)Gui, Li, and Ji], also referred to as human-in-the-loop test-time adaptation (HILTTA) throughout this study. SimATTA [Gui et al.(2024)Gui, Li, and Ji] demonstrated that adapting pre-trained models with labeled testing data could substantially improve the effectiveness and prevent catastrophic forgetting [Kemker et al.(2018)Kemker, McClure, Abitino, Hayes, and Kanan, Li and Hoiem(2017)].
@@
Figure 1: An illustration of Human-in-the-Loop Test-Time Adaptation framework. Upon collecting a batch of testing data $\{x_{i}\}_{i=1\cdots N_{b}}$, we train multiple candidate models $\theta^{*}(\omega_{m})$ using candidate hyper-parameters $\{\omega_{m}\}$. We further select candidate samples for annotation through K-Margin. Model selection is achieved by choosing the hyper-parameter minimizing validation objective. We repeat until testing data stream terminates.
Although theoretically guaranteed and offering superior results, the current implementation of the HILTTA method adheres to the conventional path of research in active learning [Sener and Savarese(2018), Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash]. In this study, we aim to explore the question: How can human-in-the-loop test-time adaptation be realized beyond an active learning perspective? To answer this question, we first revisit the pitfalls of state-of-the-art test-time adaptation [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin]. Among the multiple remaining pitfalls, we pinpoint hyper-parameter selection as the most nasty one due to the high risk of selecting inappropriate hyper-parameters by many TTA methods. Therefore, we offer a fresh perspective on human-in-the-loop test-time adaptation by synergizing active learning and model selection.
In specific, upon observing a testing sample, the model will first make an instant prediction to minimize inference latency. When a fixed number of testing samples are observed, we employ a data selection strategy to recommend a subset of testing samples for annotation. The labeled data will then be utilized as the validation set for selecting the best hyper-parameters. Finally, the model is further updated with the labeled data in a fully supervised fashion. We argue that the computation cost will not significantly increase by the introduced model selection procedure, especially compared with the time required for human labeling.
@@
To enable the novel HILTTA procedure, we identify two major challenges. First, the validation objective deserves delicate design due to the small validation set. Without seeing the whole picture of testing data, a naive model selection strategy may overfit to the labeled testing data within a local time window. Second, the data selection strategy must be well-calibrated for HILTTA. Employing the data acquisition functions solely developed for either active learning [Sener and Savarese(2018)] or active testing [Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] is suboptimal.
@@
To tackle these challenges, we first propose a hyper-parameter regularization technique by considering the deviation of model predictions from a frozen source model. We further combine the ranking scores of validation loss and the deviation as the scoring function for model selection. To accommodate the data selection strategy for HILTTA synergizing active learning and model selection, we prioritize both model uncertainty and sample diversity through a simple re-weighting. We summarize the contributions of this work as follows.
@@
  • We approach Human-In-the-Loop Test-Time Adaptation (HILTTA) by synergizing active learning and model selection.
@@
  • We demonstrate that the proposed method is compatible with state-of-the-art off-the-shelf TTA methods. The combined methods outperform dedicated active TTA method and stream-based active learning method on four test-time adaptation datasets.
@@ 2 Related Work
Test-Time Adaptation: Traditional approaches towards unsupervised domain adaptation (UDA) [Ganin and Lempitsky(2015)] learns indiscriminative features. Simultaneously training on both source and target domain data is often infeasible due to privacy issues or storage overhead. Source-free domain adaptation (SFDA) tackles this issue and adapts model weights to target domain data only [Liang et al.(2020)Liang, Hu, and Feng, Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Yang et al.(2021)Yang, Wang, Van De Weijer, Herranz, and Jui, Liang et al.(2021)Liang, Hu, Wang, He, and Feng]. In contrast, test-time adaptation (TTA) emerges in response to adapting to an unknown target domain at the inference stage [Su et al.(2022)Su, Xu, and Jia, Sun et al.(2020)Sun, Wang, Liu, Miller, Efros, and Hardt, Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Wang et al.(2022)Wang, Fink, Van Gool, and Dai, Li et al.(2023)Li, Xu, Su, and Jia]. The success of prevailing TTA methods owns to self-training on unlabeled testing data. To prevent the model from being misled by incorrect pseudo labels, a.k.a. confirmation bias [Arazo et al.(2020)Arazo, Ortego, Albert, O’Connor, and McGuinness], TTA is achieved by partial model weights update [Iwasawa and Matsuo(2021), Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell], adding regularization [Liu et al.(2021)Liu, Kothari, Van Delft, Bellot-Gurlet, Mordan, and Alahi, Su et al.(2024a)Su, Xu, and Jia] and pseudo label thresholding [Su et al.(2024b)Su, Xu, Li, and Jia, Niu et al.(2023)Niu, Wu, Zhang, Wen, Chen, Zhao, and Tan]. Nevertheless, existing approaches are mainly built upon a fully unsupervised adaptation protocol. We argue that sparse human annotation budgets might be available and exploiting annotation budget in TTA, referred to as human-in-the-loop TTA, has gained attention from the community [Gui et al.(2024)Gui, Li, and Ji]. Contrary to the traditional active learning perspective [Gui et al.(2024)Gui, Li, and Ji], we identified an unexplored opportunity to treat labeled data as the validation set for model selection and we propose a novel HILTTA paradigm by synergizing active learning and model selection.
@@
Active Learning: Active learning (AL) aims to select the most informative samples for Oracle to label for supervised training. Prevailing approaches often prioritize uncertainty [Gal and Kendall(2017)], feature diversity [Sener and Savarese(2018)] or the combination of both [Ash et al.(2020)Ash, Zhang, Krishnamurthy, Langford, and Agarwal]. The traditional AL approaches often adopt a batch-wise paradigm. When data arrive in a sequential manner stream-based AL [Saran et al.(2023)Saran, Yousefi, Krishnamurthy, Langford, and Ash] proposed to quantify the increment of the determinant as the criteria for incremental sample selection. Active learning was integrated into test-time adaptation to tackle the forgetting issue [Gui et al.(2024)Gui, Li, and Ji]. In contrast to tackling HILTTA from a pure active learning perspective, we propose to synergize active learning and model selection. Selecting samples for model selection often relies on an orthogonal objective to active learning. In the related works, active testing [Kossen et al.(2021)Kossen, Farquhar, Gal, and Rainforth, Kossen et al.(2022)Kossen, Farquhar, Gal, and Rainforth] was developed to select a subset of testing data that is representative of the whole testing data. Active testing has been demonstrated to help label-efficient model validation. Nevertheless, active learning and active testing will prioritise supervised training and model validation respectively. We believe either is not enough for for purpose of synergizing active learning and model selection for test-time adaptation.
@@ 3.1 Human-in-the-Loop Test-Time Adaptation protocol
We first provide an overview of the existing test-time adaptation (TTA) evaluation protocol. TTA aims to adapt a pre-trained model $h(x;\theta)$ to a target testing data stream ${D}=\{x_{i},\bar{y}_{i}\}_{i=1\cdots N_{t}}$ where $x_{i}$ and $\bar{y}_{i}\in\{1\cdots C\}$ denote the testing sample and the associated unknown label, respectively. Existing test-time adaptation approaches often make an instant prediction $\hat{y}=h(x;\theta)$ on each testing sample and then update model weights $\theta$ upon observing a cumulative batch of samples ${B}_{t}=\{x_{i}\}_{i=1\cdots N_{b}}$ where $t$ refers to the $t$-th minibatch. Prevailing methods enable adaptation by self-training [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell, Su et al.(2024b)Su, Xu, Li, and Jia] or distribution alignment [Su et al.(2022)Su, Xu, and Jia]. Unlike the existing assumption of totally unsupervised adaptation, additional human annotation ${B}^{l}_{t}=\{x_{j},y_{j}\}_{j=1\cdots N_{b}^{l}}$ within each batch could be introduced [Gui et al.(2024)Gui, Li, and Ji] and TTA is implemented on the combine of large unlabeled data ${B}_{t}^{u}={B}_{t}\setminus{B}_{t}^{l}$ and sparsely labeled data ${B}^{l}_{t}$. We refer to the novel evaluation protocol as Human-in-the-Loop Test-Time Adaptation (HILTTA) throughout the study.
@@ 3.2 Model Selection with Sparse Annotation
We present the details for model selection with sparse human annotation on the testing data stream. The choice of hyper-parameters is known to have a significant impact on the performance of model training and generalization. W.l.o.g., we define the task of model selection by introducing a fixed candidate pool of hyper-parameters $\Omega=\{\omega_{m}\}_{m=1\cdots N_{m}}$ with $N_{m}$ options. The choice of hyper-parameters hinges on the sensitivity of respective TTA methods, e.g. pseudo label threshold, learning rate, and loss coefficient are commonly chosen hyper-parameters for selection. We further denote the discriminative model trained upon the hyper-parameters $\omega$ as $\theta(\omega)$. The model selection can be formulated as a bi-level optimization problem as follows, where $\mathcal{L}_{tr}(\cdot;\omega)$ denotes the training loss dictated by the hyper-parameter $\omega$, $\mathcal{L}_{val}(\cdot,\cdot)$ denotes the validation loss/objective and ${D}^{l}$ refers to the labeled data.
$\begin{split}&\min_{\omega\in{\Omega}}\frac{1}{N_{t}^{l}}\sum_{x_{j},y_{j}\in{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta^{*}(\omega)),y_{j})\quad s.t.\;\theta^{*}(\omega)=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in{D}}\mathcal{L}_{tr}(h(x_{i});\omega)\end{split}$
We interpret the above bi-level optimization problem as discovering the optimal hyper-parameter $\omega$ trained upon which the model weights $\theta^{*}$ achieves the best performance on the labeled dataset. Solving the above problem through gradient-based method [Liu et al.(2018)Liu, Simonyan, and Yang] is infeasible for TTA tasks where adaptation speed is a major concern while gradient-based methods require iteratively updating between meta gradient descent $\nabla_{\omega}\mathcal{L}_{val}$ and task gradient descent $\nabla_{\theta}\mathcal{L}_{tr}$. Considering the task gradient descent only takes a few steps in a typical TTA setting, we opt for an exhaustive search for the hyper-parameters in the discretized hyper-parameter space. In specific, we enumerate the candidate models adapted with different hyper-parameters as $\{\theta^{*}_{m}\}_{m=1\cdots|\Omega|}$. The candidate with the best validation loss is chosen as the best model $\theta^{*}$, as follows.
$\begin{split}&\forall m=1\cdots N_{m},\quad\theta^{*}_{m}=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in{D}}\mathcal{L}_{tr}(h(x_{i});\omega_{m})&\theta^{*}=\arg\min_{\theta\in\{\theta^{*}_{m}\}}\sum_{x_{j},y_{j}\in{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta),y_{j})\end{split}$

### v2

@@ Abstract
Existing test-time adaptation (TTA) approaches often adapt models with the unlabeled testing data stream. A recent attempt relaxed the assumption by introducing limited human annotation, referred to as Human-In-the-Loop Test-Time Adaptation (HILTTA) in this study. The focus of existing HILTTA studies lies in selecting the most informative samples to label, a.k.a. active learning. In this work, we are motivated by a pitfall of TTA, i.e. sensitivity to hyper-parameters, and propose to approach HILTTA by synergizing active learning and model selection. Specifically, we first select samples for human annotation (active learning) and then use the labeled data to select optimal hyper-parameters (model selection). To prevent the model selection process from overfitting to local distributions, multiple regularization techniques are employed to complement the validation objective. A sample selection strategy is further tailored by considering the balance between active learning and model selection purposes. We demonstrate on 5 TTA datasets that the proposed HILTTA approach is compatible with off-the-shelf TTA methods and such combinations substantially outperform the state-of-the-art HILTTA methods. Importantly, our proposed method can always prevent choosing the worst hyper-parameters on all off-the-shelf TTA methods. The source code will be released upon publication.
**footnotetext: Equal contribution.footnotetext: Corresponding author. Email: alex.xun.xu@gmail.com
@@ 1 Introduction
Deep learning has demonstrated remarkable success in numerous application scenarios. Nevertheless, the disparity in distribution between training and testing data leads to poor generalization of deep neural networks. In many real-world scenarios, the data distribution of testing data is often unknown until the inference stage (Sun et al., 2020; Wang et al., 2020) and the distribution may experience constant shift over the course of inference (Wang et al., 2022; Yuan et al., 2023; Press et al., 2023; Lee & Chang, 2024; Su et al., 2024a). These practical challenges render traditional unsupervised domain adaptation and source-free domain adaptation paradigms ineffective. To overcome the above challenges, test-time adaptation (TTA) emerges by adapting model weights upon observing testing data at inference stage (Wang et al., 2020; Sun et al., 2020; Wang et al., 2022; Su et al., 2022; 2024a). The success of TTA is often attributed to the practice of aligning testing data distribution with source domain distribution (Su et al., 2022; Kang et al., 2023; Wang et al., 2023a) and/or self-training on pseudo labels (Jang et al., 2023; Su et al., 2024b; Marsden et al., 2024a). Despite achieving impressive results on a wide range of tasks, e.g. classification (Sun et al., 2020), image segmentation (Wang et al., 2023b) and object detection (Chen et al., 2023), the fully unsupervised TTA paradigm is still overwhelmed by remaining challenges including hyper-parameter tuning (Zhao et al., 2023), continual distribution shift (Wang et al., 2022), non-i.i.d. (Gong et al., 2022; Niu et al., 2023; Yuan et al., 2023), class imbalance (Su et al., 2024a), etc. Unfortunately, there are still no principled solutions to address these challenges without relaxing the assumptions.
@@
Figure 1: An illustration of Human-in-the-Loop Test-Time Adaptation framework. Upon collecting a batch of testing data $\{x_{i}\}_{i=1\cdots N_{b}}$, we train multiple candidate models $\theta^{*}(\omega_{m})$ using candidate hyper-parameters $\{\omega_{m}\}$. We further select candidate samples for annotation through K-Margin. Model selection is achieved by choosing the hyper-parameter minimizing validation objective. We repeat until the testing data stream concludes, following a continual TTA framework.
Recently, efforts have been made to the exploration of introducing a limited annotation budget during test-time adaptation (SimATTA) (Gui et al., 2024). In this study, we refer to this concept as human-in-the-loop test-time adaptation (HILTTA). Through a theoretical analysis, SimATTA (Gui et al., 2024) demonstrated that adapting pre-trained model with carefully selected labeled samples on the testing data stream could substantially improve the effectiveness of TTA and prevent catastrophic forgetting (Kemker et al., 2018; Li & Hoiem, 2017) upon adapting to continuously changing testing domains (Wang et al., 2022). Apart from proving the theoretical guarantees, the main focus of SimATTA (Gui et al., 2024) lies in developing a balanced incremental clustering approach to select the most informative samples on the testing data stream for annotation. Specifically, SimATTA takes into consideration the model uncertainty and feature diversity for selecting the most informative samples to annotate on the testing data stream.
Despite the theoretical guarantee and superior empirical observations, the underlying practice of utilizing label budget by existing active HILTTA methods follows the traditional line of research on active learning (Sener & Savarese, 2018; Saran et al., 2023). In this study, we aim to explore the question: How can human-in-the-loop test-time adaptation be realized beyond an active learning perspective? To answer this question, we first revisit the pitfalls of state-of-the-art test-time adaptation (Zhao et al., 2023). Among the multiple remaining pitfalls, we pinpoint hyper-parameter selection as the most nasty one due to the sensitivity to the choice of hyper-parameters by many TTA methods. Hyper-parameter selection, also referred to as model selection, aims to choose the optimal hyper-parameters for model training. Established paradigms towards model selection often rely on a held-out validation set and the model selection task is defined as discovering the hyper-parameters trained upon which the model displays optimal generalization performance on the held-out validation set (Yang & Shami, 2020). Inspired by the success of hyper-parameter optimization, we offer a fresh perspective on human-in-the-loop test-time adaptation by synergizing active learning and model selection. In specific, we present an HILTTA procedure with three steps. First, upon observing a testing sample, the model will make an instant prediction to minimize inference latency. When a fixed number of testing samples are observed, we employ a data selection strategy to recommend a subset of testing samples for human oracle to label. The labeled data will then be utilized as the validation set for selecting the best hyper-parameters. Finally, the model is further updated with the labeled data in a fully supervised fashion. With the novel HILTTA steps, we intend to make better utilization of the limited annotation budget compared to existing HILTTA approaches.
@@
To enable the novel HILTTA procedure, we identify two major challenges. First, the validation objective deserves delicate design due to the small validation set. In particular, without seeing the whole picture of testing data, a naive model selection strategy may overfit to the labeled testing data within a local time window, resulting in unstable and biased hyper-parameter selection. Second, the strategy for selecting testing samples to annotate must be well-calibrated for the task of HILTTA. Employing the data acquisition functions developed for generic active learning (Sener & Savarese, 2018) or active testing (Kossen et al., 2022) is suboptimal.
@@
To tackle these challenges, we first propose a hyper-parameter regularization technique by considering the deviation of model predictions from a frozen source model. We further combine the ranking scores of validation loss and the deviation as the scoring function for model selection. To accommodate the active learning strategy for HILTTA from a model selection perspective, we prioritize both model uncertainty and sample diversity and combine the two objectives through a simple re-weighting. Sample selection is then achieved by maximizing the coverage in the re-weighted feature space. Extensive evaluations on TTA datasets demonstrate the effectiveness of the proposed HILTTA approach.
We summarize the contributions of this work as follows.
@@
  • To the best of our knowledge, we are the first to approach Human-In-the-Loop Test-Time Adaptation (HILTTA) by synergizing active learning and model selection and propose an off-the-shelf HILTTA method.
@@
  • We demonstrate that the proposed method is compatible with state-of-the-art TTA methods. The combined methods outperform dedicated active TTA methods and stream-based active learning methods on five test-time adaptation datasets.
@@ 2.1 Test-Time Adaptation
The generalization of deep learning model is hindered by the distribution gap between source and target domain. Traditional approaches towards improving the generalization of deep learning models follow unsupervised domain adaptation (UDA) (Ganin & Lempitsky, 2015) which adapts model weights to learn indiscriminative features across source and target domains. Simultaneously training on both source and target domain data is infeasible in many real-world scenarios where access to source training data is prohibited due to privacy or storage overhead. Source-free domain adaptation (SFDA) targets this issue and adapts model weights to target domain data only (Liang et al., 2020; Liu et al., 2021; Yang et al., 2021; Liang et al., 2021). In contrast to the assumptions made in UDA and SFDA, test-time adaptation (TTA) emerges in response to the need for adapting pre-trained model to unknown target domain at inference stage (Su et al., 2022; Sun et al., 2020; Wang et al., 2020). TTA has been demonstrated to improve model’s generalization to out-of-distribution testing data and the success owns to self-training on unlabeled testing data. To prevent the model from being mislead by incorrect pseudo labels, a.k.a. confirmation bias (Arazo et al., 2020), people choose to update only a fraction of model weights, e.g. classifier weights (Iwasawa & Matsuo, 2021) and normalization layers (Wang et al., 2020), or introducing regularizations, e.g. distribution alignment (Liu et al., 2021; Su et al., 2024b) and representation regularization (Su et al., 2024a). More recent focus of test-time adaptation research lies on tackling challenges arising from real-world testing data, including continual distribution shift (Wang et al., 2022), class imbalance (Su et al., 2024a), update with small batchsize (Niu et al., 2023) and open-world testing data (Li et al., 2023; Zhou et al., 2023). Nevertheless, existing TTA approaches are mainly built upon a fully unsupervised adaptation protocol. We argue that sparse human annotation budgets might be available and exploiting annotation budget in TTA, referred to as human-in-the-loop TTA, has gained attention from the community (Gui et al., 2024). Contrary to the traditional active learning perspective (Gui et al., 2024) we identified an unexplored opportunity to treat labeled data as the validation set for model selection and we propose multiple techniques to improve the efficacy of human-in-the-loop test-time adaptation.
2.2 Active Learning
@@
Active learning (AL) aims to select the most informative samples for Oracle to label for supervised training. Prevailing approaches often prioritize uncertainty (Gal & Kendall, 2017), feature diversity (Sener & Savarese, 2018) or the combination of both (Ash et al., 2020). The traditional AL approaches often adopt a batch-wise paradigm. When data arrive in a sequential manner stream-based AL (Saran et al., 2023) proposed to quantify the increment of the determinant as the criteria for incremental sample selection. Active learning was integrated into test-time adaptation to tackle the forgetting issue (Gui et al., 2024). In contrast to tackling HILTTA from a pure active learning perspective, we propose to synergize active learning and model selection. Selecting samples for model selection often relies on an orthogonal objective to active learning. In the related works, active testing (Kossen et al., 2021; 2022) was developed to select a subset of testing data that is representative of the whole testing data. Active testing has been demonstrated to help label-efficient model validation. Nevertheless, active learning and active testing will prioritize supervised training and model validation respectively. We believe either is not enough for the purpose of synergizing active learning and model selection for test-time adaptation.
@@ 3.1 Human-in-the-Loop Test-Time Adaptation Protocol
We first provide an overview of the existing test-time adaptation (TTA) evaluation protocol. TTA aims to adapt a pre-trained model $h(x;\theta)$ to a target testing data stream $\mathbb{D}=\{x_{i},\bar{y}_{i}\}_{i=1\cdots N_{t}}$ where $x_{i}$ and $\bar{y}_{i}\in\{1\cdots C\}$ denote the testing sample and the associated unknown label, respectively. Existing test-time adaptation approaches often make an instant prediction $\hat{y}=h(x;\theta)$ on each testing sample and then update model weights $\theta$ upon observing a cumulative batch of samples $\mathbb{B}_{t}=\{x_{i}\}_{i=1\cdots N_{b}}$ where $t$ refers to the $t$-th minibatch. Prevailing methods enable adaptation by self-training (Wang et al., 2020; Su et al., 2024b) or distribution alignment (Su et al., 2022). Unlike the existing assumption of totally unsupervised adaptation, additional human annotation $\mathbb{B}^{l}_{t}=\{x_{j},y_{j}\}_{j=1\cdots N_{b}^{l}}$ within each batch could be introduced (Gui et al., 2024) and TTA is implemented on the combine of large unlabeled data $\mathbb{B}_{t}^{u}=\mathbb{B}_{t}\setminus\mathbb{B}_{t}^{l}$ and sparsely labeled data $\mathbb{B}^{l}_{t}$. We refer to the novel evaluation protocol as Human-in-the-Loop Test-Time Adaptation (HILTTA) throughout the study.
@@ 3.2 Model Selection with Sparse Annotation
We present the details for model selection with sparse human annotation on the testing data stream. The choice of hyper-parameters is known to have a significant impact on the performance of model training and generalization. W.l.o.g., we define the task of model selection by introducing a fixed candidate pool of hyper-parameters $\Omega=\{\omega_{m}\}_{m=1\cdots N_{m}}$ with $N_{m}$ options. The choice of hyper-parameters hinges on the sensitivity of respective TTA methods, e.g. pseudo label threshold, learning rate, and loss coefficient are commonly chosen hyper-parameters for selection. We further denote the discriminative model trained upon the hyper-parameters $\omega$ as $\theta(\omega)$. The model selection can be formulated as a bi-level optimization problem as follows, where $\mathcal{L}_{tr}(\cdot;\omega)$ denotes the training loss dictated by the hyper-parameter $\omega$, $\mathcal{L}_{val}(\cdot,\cdot)$ denotes the validation loss/objective and $\mathbb{D}^{l}$ refers to the labeled data.
$\begin{split}&\min_{\omega\in\Omega}\frac{1}{N_{t}^{l}}\sum_{(x_{j},y_{j})\in\mathbb{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta^{*}(\omega)),y_{j})\qquad\text{s.t.}\quad\theta^{*}(\omega)=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in\mathbb{D}}\mathcal{L}_{tr}(h(x_{i};\theta),\omega)\end{split}$ (1)
We interpret the above bi-level optimization problem as discovering the optimal hyper-parameter $\omega$ trained upon which the model weights $\theta^{*}$ achieves the best performance on the labeled dataset. Solving the above problem through gradient-based method (Liu et al., 2018) is infeasible for TTA tasks where adaptation speed is a major concern while gradient-based methods require iteratively updating between meta gradient descent $\nabla_{\omega}\mathcal{L}_{val}$ and task gradient descent $\nabla_{\theta}\mathcal{L}_{tr}$. Considering the task gradient descent only takes a few steps in a typical TTA setting, we opt for an exhaustive search for the hyper-parameters in the discretized hyper-parameter space. In specific, we enumerate the candidate models adapted with different hyper-parameters as $\{\theta^{*}_{m}\}_{m=1\cdots|\Omega|}$. The candidate with the best validation loss is chosen as the best model $\theta^{*}$, as follows.
$\begin{split}\forall m=1,\ldots,N_{m},\quad&\theta^{*}=\arg\min_{\theta\in\{\theta^{*}_{m}\}}\sum_{x_{j},y_{j}\in\mathbb{D}^{l}}\mathcal{L}_{val}(h(x_{j};\theta),y_{j})\\ \text{s.t.}\quad&\theta^{*}_{m}=\arg\min_{\theta}\frac{1}{N_{t}}\sum_{x_{i}\in\mathbb{D}}\mathcal{L}_{tr}(h(x_{i};\theta),\omega_{m})\end{split}$ (2)